OpenStudy (anonymous):

Your friend is trying to calculate the height of a nearby oak tree. You tell him that you learned how to use similar triangles in Geometry class. You tell your friend to measure his height (75 inches) and you measure the length of his shadow (48 inches). Both of you measure the length of the tree’s shadow (38 feet). How tall is the tree (in feet)? Round to the nearest hundredth. help please!!1

Directrix (directrix):

75" is to the tree's height as 48" is to 38'. Have to change inches to feet so ...

Directrix (directrix):

6.25 / t = 4 / 38 Solve for t, height of tree and ...

Directrix (directrix):

6.25/t = 2/19

Directrix (directrix):

2t = 19(6.25) t = 59.38 feet

Directrix (directrix):

Question?

OpenStudy (anonymous):

No, thank you sooo much =)